QUESTION 22QUESTION 22 Consider the information given on question 21, the Friedman test statistic is 1. 4.9 2. 5.1 3. 3.86 4. 10.61 5. 4QUESTION 21 Using the Friedman Test, determine whether there is enough evidence provided by the accompa- nying data to enable

QUESTION 22 Consider the information given on question 21, the Friedman test statistic is 1. 4.9 2. 5.1 3. 3.86 4. 10.61 5. 49
QUESTION 21 Using the Friedman Test, determine whether there is enough evidence provided by the accompa- nying data to enable to enter that at least two population locations differ (use a 0.05). Sample 1 Sample 2 Sample 3 25 19 21 23 27 25 20 29 128 23 28
Answer #1

4.9

X-r = (12/(nk(k+1 )) * (JP)-3nkti) X-,-0.2 * 324.5-60 r= 4.9